# Rational Design Heterobilayers Photocatalysts for Efficient Water Splitting Based on 2D Transition-Metal Dichalcogenide and Their Janus

### v1

@@ 3.5 Optical properties
The optical properties of $\mathrm{WSe_{2}}$-SWSe, $\mathrm{WSe_{2}}$-TeWSe, and $\mathrm{WS_{2}}$-SMoSe are calculated by using the independent particle approximation. The absorption coefficient $\alpha(\omega)$ is obtained from the calculated real $\varepsilon_{1}$ and imaginary $\varepsilon_{2}$ parts of the dielectric function as follows 42:
@@
Here, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ are scaled by the factor $h/d_{0}$ where $h$ represents the unit cell thickness and $d_{0}$ denotes the effective material thickness 14. The effective thickness, $d_{0}$, is calculated using the formula: $d_{0}=d_{\text{outer}}+r_{\text{atom1}}+r_{\text{atom2}}$, where $d_{\text{outer}}$ is the distance between the two outermost atoms of the heterostructure, as illustrated in 1. The terms $r_{\text{atom1}}$ and $r_{\text{atom2}}$ correspond to the vdW radii of the two outermost atoms 61.
The absorption spectra exhibit strong anisotropic behavior, with $\alpha(\omega)$ along the in-plane ($E_{\parallel}$) reaching up to $0.84\times 10^{6}$ cm-1 in the visible light region for $\mathrm{WS_{2}}$-SMoSe and WSe2-TeWSe, and $0.77\times 10^{6}$ cm-1 for WSe2-SWSe. These values surpass those of monolayer $\gamma$-GeSSe ($0.44\times 10^{6}$ cm-1) 1 and PdSSe ($0.24\times 10^{6}$ cm-1) 62. Notably, the $E_{\parallel}$ optical absorption differs considerably from the out-of-plane ($E_{\perp}$) absorption, especially within the visible spectrum. The combination of superior STH efficiency and strong light absorption across all materials highlights their potential for optoelectronic and water-splitting applications.
@@ 3.6.1 Gibbs free energy steps for HER
$\Delta E_{\text{H}}=E_{\text{Janus-TMDC + H}}-E_{\text{Janus-TMDC}}-\frac{1}{2}E_{\text{H}},$ (16)
where $E_{\text{Janus-TMDC + H}}$ is the total energy of the heterostructure of Janus monolayer and TMDC material with an adsorbed hydrogen atom, $E_{\text{Janus-TMDC}}$ is the energy of the bare vdW layers, and $\frac{1}{2}E_{\text{H}}$ represents the energy of a hydrogen molecule in its standard state. Building upon this, the potential provided by photogenerated electrons for HER ($U_{e}$) can be calculated using the equation $U_{e}=V_{\text{CB}}-4.027$ for pH = 7 14, 63, where $U_{e}$ corresponds to the electron potential associated with the CBM, potentially matching the electron affinity $\chi(\text{H}_{2})$ under illuminated conditions. In a dark environment, where no photogenerated electrons are available, $U_{e}$ becomes zero, indicating the absence of a driving force for HER. Based on the calculations in the previous section, the $U_{e}$ values for $\text{WS}_{2}$-SMoSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe under the illumination were found to be 0.10, 0.12, and 0.09 eV, respectively.
We further evaluated the HER activity at the reduction sites of $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and again $\text{WS}_{2}$-SMoSe, with the corresponding $\Delta G_{\text{H}}$ values presented in Figure 5. The calculated free energy changes without an external potential are 2.10, 2.18, and 2.27 eV for these materials, respectively. These high positive values of $\Delta G_{\text{H}}$ indicate that spontaneous hydrogen adsorption does not occur easily, implying limited catalytic efficiency under standard conditions. Nevertheless, this behavior is similar to other well-known materials such as $\text{MoS}_{2}$, $\text{MoSe}_{2}$, $\text{WSe}_{2}$, and the Janus MoSSe materials 64, 65, 66. On the other hand, calculated free energy changes with an external potential are the $\Delta G_{\text{H}}$ values to 2.01, 2.06, and 2.17 eV for $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe, respectively. These reduction values suggest that suitable irradiation can enhance the HER performance by decreasing the energy barrier for hydrogen adsorption. Although the $\Delta G_{\text{H}}$ values remain positive, indicating that the hydrogen evolution reaction is still non-spontaneous, the decreased values under light exposure or applied potential could potentially improve the catalytic performance by making hydrogen releasing more favorable.
@@ 3.6.2 Diffusion of H atom on the HER sites
To investigate H atom diffusion on the HER-active surfaces, we calculate the optimized migration pathways using a minimal energy approach 45. The pathways and the corresponding energy profiles are shown in Figure 6, where S1 $\rightarrow$ S2 represents H atom migration between adjacent hex-atomic-rings, and S2 $\rightarrow$ S3 denotes migration within a single hex-atomic-ring. The migration pathways are optimized by the nudged elastic band (NEB) method 45 with a convergence criterion of 0.05 eV/Å for NEB calculations with 13 intermediate images. For WS2-SMoSe, H migration on the SMoSe surface showed a lower energy barrier for inter-ring migration (0.199 eV) than intra-ring migration (0.216 eV). In the case of WSe2-TeWSe, the WSe2 surface exhibited easier migration between hex-atomic-rings (0.149 eV) than within a ring (0.286 eV). For WS2-SMoSe, the SMoSe surface favored intra-ring migration (0.200 eV) over inter-ring movement (0.224 eV), with this structure exhibiting a lower energy barrier compared to the other two. Compared to previously reported values 45, the migration energy barriers of these structures are lower than those of monolayer MoSSe, which were observed at 0.32 eV for inter-ring and 0.08 eV for intra-ring migration on Se-site. This reduction suggests that these WS2-SMoSe, WSe2-TeWSe, and WS2-SMoSe surfaces provide more favorable diffusion pathways, which could enhance HER activity on these materials.
@@ 4 Conclusions
The findings in this study underscore the potential of Janus-TMDC heterobilayers as highly efficient photocatalysts for STH conversion. By leveraging intrinsic electric fields generated through Janus asymmetry, we achieve enhanced charge separation and minimized electron-hole recombination in Z-scheme configurations such as WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe. These Z-scheme structures not only enable spatially separated sites for HER and OER but also provide favorable band alignment for water splitting, with WS2-SMoSe achieving the highest STH efficiency of 33.24%. In exploring the relationship between carrier mobility and photocatalytic capabilities, we use the Fröhlich interaction model to account for LO phonon scattering, revealing that mobility in Janus-TMDC heterobilayers is significantly affected by their intrinsic polarization. Additionally, HER analysis shows that external potential effectively lowers reaction barriers, enhancing HER efficiency under illumination. Hydrogen diffusion also indicates relatively low energy barriers for H atom migration on HER-active surfaces, suggesting efficient diffusion on the WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe surfaces. Overall, this work provides a comprehensive framework for selecting and designing the optimized Janus-TMDC heterobilayers for clean energy applications, presenting a promising pathway toward advancing efficient photocatalytic systems.
@@
Supporting Information: Table of structural parameters and band Gap; table of Born charges, Phonon frequency, and dielectric properties; figure of HSE06-calculated band structure; figure of electrostatic potential along vacuum axis ($z$); figure of band alignments for water splitting at pH 0, 2, and 7; figure of PBE band structure for the effective mass calculation; figure of total energy-strain relationship in zigzag and armchair directions; figure of CBM and VBM energies under uniaxial strain; figure of Phonon dispersion and DOS.

### v2

@@ 3.7 Optical properties
The optical properties of $\mathrm{WSe_{2}}$-SWSe, $\mathrm{WSe_{2}}$-TeWSe, and $\mathrm{WS_{2}}$-SMoSe are calculated by using the independent particle approximation. The absorption coefficient $\alpha(\omega)$ is obtained from the calculated real $\varepsilon_{1}$ and imaginary $\varepsilon_{2}$ parts of the dielectric function as follows 39:
@@
Here, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ are scaled by the factor $h/d_{0}$ where $h$ represents the unit cell thickness and $d_{0}$ denotes the effective material thickness 8. The effective thickness, $d_{0}$, is calculated using the formula: $d_{0}=d_{\text{outer}}+r_{\text{atom1}}+r_{\text{atom2}}$, where $d_{\text{outer}}$ is the distance between the two outermost atoms of the heterostructure, as illustrated in Figure 1. The terms $r_{\text{atom1}}$ and $r_{\text{atom2}}$ correspond to the vdW radii of the two outermost atoms 62.
In our study, we observed that the optical absorption spectra of TMDC/Janus heterobilayers exhibit strong anisotropic behavior. The in-plane absorption coefficient $\alpha(\omega)$ ($E_{\parallel}$) reaches up to $0.84\times 10^{6}$ cm-1 in the visible light region for $\mathrm{WS_{2}}$-SMoSe and WSe2-TeWSe, and $0.77\times 10^{6}$ cm-1 for WSe2-SWSe. These values significantly surpass those of monolayer Janus materials such as $\gamma$-GeSSe ($0.44\times 10^{6}$ cm-1) 1 and PdSSe ($0.24\times 10^{6}$ cm-1) 63, as well as monolayers of MoSSe, MoSeTe, MoSTe, WSSe, WSeTe, and WSTe 64. This substantial enhancement and broadening of absorption in TMDC/Janus heterobilayers can be attributed to strong light–matter coupling 65, 66, which is more pronounced in multilayered structures.
Notably, the in-plane optical absorption ($E_{\parallel}$) differs considerably from the out-of-plane absorption ($E_{\perp}$), emphasizing the anisotropic optical properties of these heterobilayers. The combination of superior STH efficiency and strong light absorption across all materials underscores their potential for optoelectronic and water-splitting applications. This highlights the role of the Z-scheme configuration for narrow band gap semiconductors, which creates favorable conditions for the water-splitting process by extending the intrinsic band gap while retaining the efficient light-harvesting properties of small band gap materials.
@@ 3.8.1 Gibbs free energy steps for HER
$\Delta E_{\text{H}}=E_{\text{Janus-TMDC + H}}-E_{\text{Janus-TMDC}}-\frac{1}{2}E_{\mathrm{H_{2}}},$ (18)
where $E_{\text{Janus-TMDC + H}}$ is the total energy of the heterostructure of Janus monolayer and TMDC material with an adsorbed hydrogen atom, $E_{\text{Janus-TMDC}}$ is the energy of the bare vdW layers, and $\frac{1}{2}E_{\mathrm{H_{2}}}$ represents the energy of a hydrogen molecule in its standard state. Building upon this, the potential provided by photogenerated electrons for HER ($U_{e}$) can be calculated using the equation $U_{e}=V_{\text{CB}}-E_{\text{H}^{+}/\text{H}_{2}}$ 8, 44, where $U_{e}$ corresponds to the electron potential associated with the CBM, potentially matching the electron affinity $\chi(\text{H}_{2})$ under illuminated conditions. In a dark environment, where no photogenerated electrons are available, $U_{e}$ becomes zero, indicating the absence of a driving force for HER. Based on the calculations in the previous section, the $U_{e}$ values for $\text{WS}_{2}$-SMoSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe under the illumination were found to be 0.10, 0.12, and 0.09 eV, respectively, at pH = 7. At pH = 0, the corresponding values were 0.52 eV, 0.53 eV, and 0.51 eV, respectively.
We further evaluated the HER activity at the reduction sites of $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and again $\text{WS}_{2}$-SMoSe, with the corresponding $\Delta G_{\text{H}}$ values presented in Figure 6. The calculated free energy changes without an external potential are 1.60, 1.52, and 1.43 eV for these materials, respectively. These high positive values of $\Delta G_{\text{H}}$ indicate that spontaneous hydrogen adsorption does not occur easily, implying limited catalytic efficiency under standard conditions. Nevertheless, this behavior is similar to other well-known materials such as $\text{MoS}_{2}$, $\text{MoSe}_{2}$, $\text{WSe}_{2}$, and the Janus MoSSe materials 67, 68, 69. On the other hand, calculated free energy changes with an external potential are the $\Delta G_{\text{H}}$ values to 1.50, 1.40, and 1.34 eV for $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe, respectively, at pH = 7. At pH = 0, the corresponding values were 1.08 eV, 0.99 eV, and 0.92 eV, respectively. These reduction values suggest that suitable irradiation can enhance the HER performance by decreasing the energy barrier for hydrogen adsorption. Although the $\Delta G_{\text{H}}$ values remain positive, indicating that the hydrogen evolution reaction is still non-spontaneous, the decreased values under light exposure or applied potential could potentially improve the catalytic performance by making hydrogen releasing more favorable.
@@ 3.8.2 Diffusion of H atom on the HER sites
To investigate H atom diffusion on the HER-active surfaces, we calculate the optimized migration pathways using a minimal energy approach 70. The pathways and the corresponding energy profiles are shown in Figure 7, where S1 $\rightarrow$ S2 represents H atom migration between adjacent hex-atomic-rings, and S2 $\rightarrow$ S3 denotes migration within a single hex-atomic-ring. The migration pathways are optimized by the nudged elastic band (NEB) method 70 with a convergence criterion of 0.05 eV/Å for NEB calculations with 13 intermediate images. For WS2-SMoSe, H migration on the SMoSe surface showed a lower energy barrier for inter-ring migration (0.199 eV) than intra-ring migration (0.216 eV). In the case of WSe2-TeWSe, the WSe2 surface exhibited easier migration between hex-atomic-rings (0.149 eV) than within a ring (0.286 eV). For WS2-SMoSe, the SMoSe surface favored intra-ring migration (0.200 eV) over inter-ring movement (0.224 eV), with this structure exhibiting a lower energy barrier compared to the other two. The preference of the H atom for the Janus site over the TMDC site can be attributed to the smaller electronegativity difference between the H atom and the S atom (on the Janus site) compared to the difference between the H atom and the Se atom (on the TMDC site). This distinction highlights the potential of designing Janus/TMDC heterobilayers with minimal interaction between the H atom and chalcogenide atoms on the reduction site, thereby reducing the interaction between the H atom and the material layers. Such design could lead to more efficient catalytic performance.
Compared to previously reported values 70, the migration energy barriers of these structures are lower than those of monolayer MoSSe, which were observed at 0.32 eV for inter-ring and 0.08 eV for intra-ring migration on the Se site. This reduction suggests that the WS2-SMoSe, WSe2-TeWSe, and WS2-SMoSe surfaces provide more favorable diffusion pathways. Furthermore, thanks to the Z-scheme heterostructure alignment, the electric field was modified, which contributed to reducing the interaction between the H atom and the photocatalyst system, as observed in the 3 calculated samples 71. This modification plays a role in enhancing the efficiency of the HER.
@@ 4 Conclusions
The findings in this study underscore the potential of Janus-TMDC heterobilayers as highly efficient photocatalysts for STH conversion. By leveraging intrinsic electric fields generated through Janus asymmetry, we achieve enhanced charge separation and minimized electron-hole recombination in Z-scheme configurations such as WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe. These Z-scheme structures not only enable spatially separated sites for HER and OER but also provide favorable band alignment for water splitting, with WS2-SMoSe achieving the highest STH efficiency of 16.62%. Our framework highlights the critical role of electronegativity differences and interfacial atomic configurations in determining band alignment types. Materials exhibiting significant internal electric fields and aligned synthetic and internal electric field directions favor Type-II configurations, while opposing field directions lead to Z-scheme alignments. This predictive framework aids in the rational design of heterobilayers with tailored electronic properties, advancing their application in photocatalysis and beyond. In exploring the relationship between carrier mobility and photocatalytic capabilities, we use the Fröhlich interaction model to account for LO phonon scattering, revealing that mobility in Janus-TMDC heterobilayers is significantly affected by their intrinsic polarization. Additionally, HER analysis shows that external potential effectively lowers reaction barriers, enhancing HER efficiency under illumination. Hydrogen diffusion also indicates relatively low energy barriers for H atom migration on HER-active surfaces, suggesting efficient diffusion on the WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe surfaces. Overall, this work not only establishes a comprehensive framework for evaluating and predicting the performance of Janus-TMDC heterobilayers but also paves the way for innovative design strategies. By integrating insights from band alignment, intrinsic electric fields, and carrier mobility, we present a holistic approach to optimizing these heterobilayers for clean energy applications. The findings provide a robust foundation for advancing next-generation photocatalytic systems, offering transformative potential for hydrogen production, light harvesting, and other energy-related technologies.
@@
Supporting Information: Table of structural parameters and band gap; table of Born charges, Phonon frequency, and dielectric properties; figure of HSE06-calculated band structure; figure of electrostatic potential along vacuum axis ($z$); figure of electrostatic potential differences across interface for different alignment types; figure of band alignments for water splitting at pH 0, 2, and 7; figure of PBE band structure for the effective mass calculation; figure of total energy-strain relationship in zigzag and armchair directions; figure of CBM and VBM energies under uniaxial strain; figure of Phonon dispersion and DOS.